cmhardw
Premium Member
Totally, like, sometimes I get 4-5 PLL skips in a row.
Other times I get like, 20 PLL skips in a row.
yeah, the probability is just 1/14016833953562607293918185758734155776 .
You should play to some lotery.
(ZB method)
Yes, but we are using terminology incorrectly here
PLL skip implies that you "skipped" the step as per chance. As in you had some chance to skip this step, and you got lucky and this occurred.
The chance of a "PLL" skip for ZBLL is actually the chance of a full LL skip. The reason is that PLL is not a step in the ZB method but a special case of the ZB method.
For example, Chris has a 1/27 chance to have a fully oriented last layer after solving ZBF2L. This means that Chris will do a standard PLL alg when this happens, most of the time. Now, given that Chris has ended up with an oriented last layer after ZBF2L (1/27 chance to do so), he then has a 1/72 chance to skip the resulting PLL case.
So Chris, the chance that you have a PLL skip is 1/27 * 1/72 = 1/1944 This of course assumes you don't do anything extra crazy to partially influence the orientation of corners during ZBF2L.
Remember that PLL is not a step in your method, it is a special sub-case of your method, the one that occurs 1/27 of the time when the last layer is oriented after ZBF2L. Using terminology in this way can lead to confusing claims of "I skipped PLL 20 times in a row" and counter-claims of "This is so completely improbable as to be effectively impossible".
Chris